The probability that person A will pass Finite Mathematics is 6/11 and the probability that  person B will pass is 4/7 . Assume the events are independent. Find the probability that neither will pass.
Answer #1

P(A Pass) = 6/11 , P(B pass) = 4/7

Since Aand B are independent,

P(A pass and B pass) = P(A pass) * P(B pass)

= 6/11 * 4/7

= 24/77

P(A pass OR B pass) = P(A Pass) + P(B pass) - P(A pass and B pass)

= 6/11 + 4/7 - 24/77

= 62/77

Therefore,

P(Neither will pass) = 1 - P(A pass OR B pass)

= 1 - 62/77

= 15/77 ( OR 0.1948)